# Focused Transformer: Contrastive Training for Context Scaling

### v1

@@ 5.3 FoT fine-tuning and context length extrapolation
As a base model, we use a standard transformer model pre-trained for $100k$ steps with context of $1K$ tokens using the standard objective and fine-tune with the FoT objective (i.e. crossbatch). The data used for both fine-tuning and pre-training is the C4 dataset Raffel et al. (2019a) (we omit documents shorter than $2K$ tokens). The fine-tuning phase takes $10k$ steps. We use the crossbatch dimension $d=128$ and local context of $1K$ tokens (context is $2K$ during training). We evaluate models in a zero-shot way on $4$ language modeling datasets, which require long context: arXiv, PG-19, GitHub and Isabelle, see Section 5.1 and Appendix C for details.
@@
In Table 4, we observe that FoT enjoys steady perplexity gains up to $64K$ tokens, although it was fine-tuned only with the $2K$ total differentiable context length. We compare the model perplexity to the following baselines: Memorizing Transformer (MT) (Wu et al., 2022) fine-tuned with the local context of $1K$ and memory size of $16K$, and Transformer-XL (Dai et al., 2019) fine-tuned with both local context and window length of $1K$. To ensure a fair comparison, all three models are fine-tuned from the same base checkpoint. When evaluated with a context of $2K$, our method achieves results on par with the Transformer-XL baseline, which has access to the previous context in all layers, unlike MT and FoT. Compared to the MT baseline, we achieve better scaling when evaluated with $64K$ context length and significantly better perplexity values. Unlike MT, our method does not require training on long sequences, which is reflected by the lower perplexities of FoT when evaluated in the zero-shot setting. For more details, see Appendix E.
@@ 5.4 Handling distractions in language modeling tasks
In this section, we measure how handling distractions in the multi-document setting helps in language modeling.
We pick the PG-19 dataset (Rae et al., 2019) and measure the perplexity of the next token prediction (language modeling task) when varying the size of multi-doc memory (in this case consisting of books). Intuitively, the memory tokens corresponding to the current book might be beneficial (which is also confirmed in (Wu et al., 2022)), while the ones from the other books are unlikely to be useful and thus are distractions.
@@
Figure 5: Perplexity in multi-doc setting. FoT was trained with local context of size $512$ and different $d$. FoT $2$->$64$ was initially trained with $d=2$ and then switched to $d=64$. Single-doc MT was trained with a memory size of $16K$. As we increase the memory size, the number of distractions (irrelevant keys) increases, making the task harder. Single-doc MT evaluated in single-doc setting is a soft lower bound since it is not exposed to distractions.
@@ 5.5 Context length extrapolation in single-doc
The primary motivation behind FoT is to improve the multi-doc setting performance by handling distractions. Interestingly, our method also helps to extrapolate to longer contexts, even when evaluated in the single-doc setting.
@@ 5.6 Ablations and design choices
5.6.1 Impact of differentiable keys and values
Context Length FoT d=1 MT
512 14.18 14.68
1024 14.17 14.46
2048 14.11 14.43
5.6.2 Importance of negatives
5.6.3 Relation to Memorizing Transformer
Memorizing Transformer Wu et al. (2022) is closely related to our method. The two key differences are 1) the training protocol and 2) how the memory is integrated into the model. In this section, we provide additional insights into these differences.
Memory integration FoT uses a simple memory integration approach where the $(key,value)$ pairs retrieved by kNN lookup are treated the same way as the local context. In contrast, MT uses a gating mechanism, a weighted average of the memory, and local values; see details in Appendix B.2. We evaluated both approaches and found no difference in performance between these two memory integration methods. However, we decided to use our approach because it does not require any architectural changes (and thus makes fine-tuning existing models easy). For these reasons, we recommend using it. We speculate that the reason why the gating is not needed in FoT is another benefit of the fact that the crossbatch training backpropagates through the $(key,value)$ pairs from the previous context $C_{prev}$ in contrast to MT that cannot backpropagate there and needs to rely on local context when computing gradients for keys and values. Another reason might be the fact that $C_{prev}$ is embedded for each batch, and thus staleness (see (Wu et al., 2022, Section 3.2)) is avoided.
@@ 6 Limitations and future work
Scaling up memory This is by far the most important future research direction. The challenges start from purely engineering, storing more than $16$M $(key,value)$ pairs will require a distributed multi-node system. In our experiments, we use the exact kNN search, which is not scalable to large memory. Using approximate kNN search will require a lot of engineering effort, as well as careful evaluation of the impact of the approximation on the model performance.
Scaling up crossbatch We observed that increasing $d$ is beneficial. In our experiments, we used $d=64$ or $d=128$, which is the maximum value that fits into the memory of a single TPUv3/TPUv2 machine, see also Appendix H. In future work, we want to further increase $d$ as well as test on devices with bigger memory or utilize multi-node training.
@@ A.3 LLaMA fine-tuning dataset
All subsets apart from python are taken directly from RedPajama. For the python subset, we gather Python source code from The Stack and, to obtain long documents for training, concatenate files that are in the same subdirectory in random order, using a similar procedure as for the GitHub dataset in Section G. Additionally, we filter out short documents for some subsets of the original RedPajama, namely shorter than the Min. doc. length column indicates.
@@ B.1 Transformer models
We use the transformer architecture introduced in [Vaswani et al., 2017] with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform layer normalization before the input of both the attention and feed-forward modules. Additionally, we use Rotary Position Embedding [Su et al., 2021], normalize keys and queries [Henry et al., 2020], and introduce a learnable temperature parameter for each attention head.
The hyperparameters for each model size can be found in Appendix C. For training the models on PG-19, we use the standard T5 tokenizer with $32$k vocabulary [Raffel et al., 2019b]. The larger models in Section 5.3 are trained with a custom SentencePiece tokenizer [Kudo and Richardson, 2018] with $64$k vocabulary size.
@@ B.2 Memory attention layer
Memory attention layer $\ell$ is one of the transformer layers, which has access to the external memory $M$. The memory stores $(key,value)$ pairs. For each query $q$ in $\ell$, we retrieve the $k$ most matching entries from $M$ and use them to compute the attention value. More precisely, we use the kNN algorithm to pick $M_{top}:=\{(key_{1},value_{1}),\ldots,(key_{k},value_{k})\}\subset M$ such that $\{\langle q,key_{i}\rangle\}_{i=1,\ldots,k}$ are the top $k$ inner products in $M$. These are merged with the part of the local context before $q$ denoted as $C_{<q}$ and used to compute the attention value using the standard Transformer formula:
@@ B.5 Memorizing Transformer
We speculate that there may be benefits in blending these two approaches. One can, for example, argue that MT is better at providing ’hard’ negatives for the model. We provide a proof-of-concept experiment in Section 5.6.3, otherwise leave this for future work.
@@
Appendix C Hyperparameters
Table 9 shows hyperparameters used in our experiments. We used context length $512$ unless stated otherwise. In Section 5.2, Section 5.4, Section 5.6, we use the total batch size of $32K$ tokens. In Section 5.3 and Section 5.5, the total batch size is $128K$ tokens.
@@ Appendix C Hyperparameters
#Params $37M$ $184M$ $600M$ $1.2B$
Max learning rate $0.02$ $0.01$ $0.005$ $0.005$
Min learning rate $0.01$ $0.0005$ $0.0005$ $0.0005$
Embedding dim $512$ $1024$ $2048$ $2048$
Head dim $64$ $128$ $128$ $128$
#Heads $8$ $8$ $16$ $32$
FeedForward dim $2048$ $4096$ $8192$ $16384$
Local context length $256$/$512$ $512$ $2048$ $2048$
Batch size $32K$/$64K$ $32K$ $256K$ $256K$
#Number of training steps $5k$ $500k$ $150k$ $200k$
@@
For the dictionary lookup task (Section 5.2) we checked the following hyperparameters:
@@
C.1 Schedule of $d$
In Sections 5.4, 5.5 and 5.6 for models with $d\in\{1,2,4,8\}$ we used constant schedule, and for models with $d=64$ we trained with $d=2$ for $450k$ steps and switched to $d=64$ for the final $50k$ steps. In Section 5.2 we trained with $d=1$ until model reached $98\%$ accuracy and then we swtiched to $d=128$. For the $184M$ model in Section 5.3, we randomly sampled $d$ from $\{2,128\}$ in each training step, and the $600M$ and $1.2B$ models just used $d=2$.
Appendix D Dictionary lookup task
@@
During training, we use documents comprising $512$ tokens. The first half of each document consists of definitions, whereas the second one consists of questions. In evaluation, we use longer documents but make only the last $256$ tokens correspond to questions. That is, as the context gets bigger (the token axis on Figure 4), the number of definitions increases, but the number of queries remains the same.
Appendix E FoT fine-tuning
For comparison in Table 4, our $184M$ model is pre-trained for $100k$ steps with a total batch size of 128 ($128K$ tokens per step, with $1024$ local context). Then we fine-tune both FoT and baselines for additional $10k$ steps with the same batch size. When fine-tuning FoT, we randomly sample $d$ from $\{2,128\}$ in each training step to prevent the model from overfitting to a large additional context length during training.
Appendix F Code
In Listing 1, we show the FoTs attention code (i.e., the code for the memory attention layers and crossbatch training), see Section 3, Appendix B.2, Appendix B.3. We note that the changes to the code are small; they are localized to the memory layer (the other layers follow the standard transformer protocol) and do not require any new trainable parameters.
Listing 1: Memory attention: Let $\ell$ be a memory attention layer. During the training, we make $\ell$ attend to the $(key,value)$ pairs from the local context, previous local context, and $d-1$ contexts coming from other documents. During the inference, queries from $\ell$ attend to the local context and $k$ nearest neighbors retrieved from memory. For simplicity, we provide the code for one head and assume that the crossbatch dimension $d$ is equal to the batch size.

### v2

@@ 5.2 FoT fine-tuning and context length extrapolation
As a base model, we use a standard transformer model pre-trained for $100k$ steps with context of $1K$ tokens using the standard objective and fine-tune with the FoT objective (i.e. crossbatch). The data used for both fine-tuning and pre-training is the C4 dataset Raffel et al. (2019a) (we omit documents shorter than $2K$ tokens). The fine-tuning phase takes $10k$ steps. We use the crossbatch dimension $d=128$ and local context of $1K$ tokens (context is $2K$ during training). We evaluate models in a zero-shot way on $4$ language modeling datasets, which require long context: arXiv, PG-19, GitHub and Isabelle, see Section 5.1 and Appendix E for details.
@@
In Table 3, we observe that FoT enjoys steady perplexity gains up to $64K$ tokens, although it was fine-tuned only with the $2K$ total differentiable context length. We compare the model perplexity to the following baselines: Memorizing Transformer (MT) (Wu et al., 2022) fine-tuned with the local context of $1K$ and memory size of $16K$, and Transformer-XL (Dai et al., 2019) fine-tuned with both local context and window length of $1K$. To ensure a fair comparison, all three models are fine-tuned from the same base checkpoint. When evaluated with a context of $2K$, our method achieves results on par with the Transformer-XL baseline, which has access to the previous context in all layers, unlike MT and FoT. Compared to the MT baseline, we achieve better scaling when evaluated with $64K$ context length and significantly better perplexity values. Unlike MT, our method does not require training on long sequences, which is reflected by the lower perplexities of FoT when evaluated in the zero-shot setting. For more details, see Appendix G.
We also confirm the context extrapolation abilities using a synthetic dictionary lookup task. In this task, the model is first provided with $k_{i}:v_{i}$ mappings and then asked what value is associated with a particular key. We train $37$M parameter models using documents of length $512$. Figure 10 shows that FoT, after $5$k steps of training, can effectively utilize memory consisting of $16$M tokens achieving accuracy above $92\%$. Details can be found in Appendix F.
@@ 5.3 Handling distractions in language modeling tasks
In this section, we measure how handling distractions in the multi-document setting helps in language modeling. We pick the PG-19 dataset (Rae et al., 2019) and measure the perplexity of the next token prediction (language modeling task) when varying the size of multi-doc memory (in this case consisting of books). Intuitively, the memory tokens corresponding to the current book might be beneficial (which is also confirmed in (Wu et al., 2022)), while the ones from the other books are unlikely to be useful and thus are distractions.
@@ 5.4 Context length extrapolation in single-doc
The original motivation behind FoT is to improve the multi-doc setting performance by handling distractions. Interestingly, our method also helps to extrapolate to longer contexts, even when evaluated in the single-doc setting.
@@ 5.5 Ablations and design choices
In Appendix C we present ablations on our design choices. In particular, we note the importance of differentiability and the inclusion of negatives. We also discuss the relation to Memorizing Transformer. We note that due to the limited resources we have followed the Memorizing Transformer in the choice of memory layers.
@@ 6 Limitations and future work
Scaling up context This is by far the most important future research direction. The challenges start from purely engineering, storing more than $16$M $(key,value)$ pairs will require a distributed multi-node system. In our experiments, we use the exact kNN search, which is not scalable to large memory. Using approximate kNN search will require a lot of engineering effort, as well as careful evaluation of the impact of the approximation on the model performance.
Scaling up crossbatch We observed that increasing $d$ is beneficial. In our experiments, we used $d=64$ or $d=128$, which is the maximum value that fits into the memory of a single TPUv3/TPUv2 machine, see also Appendix I. In future work, we want to further increase $d$ as well as test on devices with bigger memory or utilize multi-node training. We also note that crossbatch increases the training cost, but only in a subset of layers.
@@
Listing 1: Possible implementation of cross-batch. To simplify the code we assume that each document occupies two consecutive elements of the batch. A more detailed version is in Appendix L.
# keys from other contexts will be encoded as if they
# were at the beginning of the local context
pkey_fst = pos_encode_as_first(xk=key)
# local context keys encoded in the standard way
pquery, pkey = pos_encode(xq=query, xk=key)
# for each element of the batch we calculate indices of
# the batch that will be used in cross-batch
cross_batch_rel_ids = jnp.arange(0, -num_attentions, -1)
                      .reshape(1, -1)
batch_ids = jnp.arange(0, batch_size).reshape(-1, 1)
cross_batch_selector = cross_batch_rel_ids + batch_ids
# here we want other contexts
cross_batch_keys = pkey_fst[cross_batch_selector[:, 1:]]
# here we concatenate local context with other contexts
attention_keys =
    jnp.concatenate([pkey[:, None], cross_batch_keys], axis=1)
cb_attn_weights =
    jnp.einsum("bqhd,bckhd->bhqck",
    pquery, attention_keys, precision=precision)
@@ A.3 LLaMA fine-tuning dataset
All subsets apart from python are taken directly from RedPajama. For the python subset, we gather Python source code from The Stack and, to obtain long documents for training, concatenate files that are in the same subdirectory in random order, using a similar procedure as for the GitHub dataset in Section H. Additionally, we filter out short documents for some subsets of the original RedPajama, namely shorter than the Min. doc. length column indicates.
In case one document is too short to span across several contexts for crossbatch, then we concatenate it with the next document from the dataset.
@@ A.4 Language Model Evaluation Harness
A.5 Question answering over research papers
We evaluate the context utilization of our model on the validation set of Qasper [Dasigi et al., 2021] from SCROLLS [Shaham et al., 2022]. Details are in the Table 6.
Table 6: Zero-shot performance with different context lengths on the validation subset of Qasper [Dasigi et al., 2021]. We use the implementation from Language Model Evaluation Harness [Gao et al., 2021a]. In the Harness implementation, yes/no questions are evaluated separately from open questions. Observe that LongLLaMA 3B benefits from the extended context.
Context length OpenLLaMA 3B LongLLaMA 3B LLaMA 7B LongChat 7B
2K 18.7 18.7 18.7 19.4
4K - 20.7 - 21.2
6K - 23.2 - 25.0
8K - 26.6 - 28.8
@@ Appendix B Architecture
This section describes the architecture and crossbatch details for non-LLaMA-based models presented in this paper. The main differences are that for LLaMA-based models (LongLLaMA) we maintain the positional encodings (with a slight modification detailed in A.2), do not introduce the attention temperature parameter, and replace kNN with full dense attention.
@@ B.1 Transformer models
For non-LLaMA-based models we use the transformer architecture introduced in [Vaswani et al., 2017] with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform layer normalization before the input of both the attention and feed-forward modules. Additionally, we use Rotary Position Embedding [Su et al., 2021], normalize keys and queries [Henry et al., 2020], and introduce a learnable temperature parameter for each attention head.
The hyperparameters for each model size can be found in Appendix E. For training the models on PG-19, we use the standard T5 tokenizer with $32$k vocabulary [Raffel et al., 2019b]. The larger models in Section 5.2 are trained with a custom SentencePiece tokenizer [Kudo and Richardson, 2018] with $64$k vocabulary size.
@@ B.2 Memory attention layer
Memory attention layer $\ell$ is one of the transformer layers, which has access to the additional context $M$. The memory stores $(key,value)$ pairs. For each query $q$ in $\ell$, we retrieve the $k$ most matching entries from $M$ and use them to compute the attention value. More precisely, we use the kNN algorithm to pick $M_{top}:=\{(key_{1},value_{1}),\ldots,(key_{k},value_{k})\}\subset M$ such that $\{\langle q,key_{i}\rangle\}_{i=1,\ldots,k}$ are the top $k$ inner products in $M$. These are merged with the part of the local context before $q$ denoted as $C_{<q}$ and used to compute the attention value using the standard Transformer formula:
@@
We do not use the $\tau$ parameter for LongLLaMAs as their architecture does not normalize keys and queries. For LongLLaMAs, we also replace the kNN search with dense attention and retain positional encodings (see Appendix A.2).
@@ B.5 Memorizing Transformer
We speculate that there may be benefits in blending these two approaches. One can, for example, argue that MT is better at providing ’hard’ negatives for the model. We provide a proof-of-concept experiment in Appendix C.3, and leave this for future work.
@@
Appendix C Ablations
C.1 Impact of differentiable keys and values
Context Length FoT d=1 MT
512 14.18 14.68
1024 14.17 14.46
2048 14.11 14.43
C.2 Importance of negatives
C.3 Relation to Memorizing Transformer
Memorizing Transformer Wu et al. [2022] is closely related to our method. The two key differences are 1) the training protocol and 2) how the memory is integrated into the model. In this section, we provide additional insights into these differences.
Memory integration FoT uses a simple memory integration approach where the $(key,value)$ pairs retrieved by kNN lookup are treated the same way as the local context. In contrast, MT uses a gating mechanism, a weighted average of the memory, and local values; see details in Appendix B.2. We evaluated both approaches and found no difference in performance between these two memory integration methods. However, we decided to use our approach because it does not require any architectural changes (and thus makes fine-tuning existing models easy). For these reasons, we recommend using it. We speculate that the reason why the gating is not needed in FoT is another benefit of the fact that the crossbatch training backpropagates through the $(key,value)$ pairs from the previous context $C_{prev}$ in contrast to MT that cannot backpropagate there and needs to rely on local context when computing gradients for keys and values. Another reason might be the fact that $C_{prev}$ is embedded for each batch, and thus staleness (see [Wu et al., 2022, Section 3.2]) is avoided.
Appendix D Additional figures
Figure 8: Perplexity in the multi-doc setting. FoT was trained with local context of size $512$ and different $d$. FoT $2$->$64$ started with $d=2$ and then switched to $d=64$. Single-doc MT was trained with a memory size of $16K$. As we increase the memory size, the number of distractions (irrelevant keys) increases, making the task harder. Single-doc MT evaluated in the single-doc setting is a soft lower bound since it lacks distractions.
Figure 8: Perplexity in the multi-doc setting. FoT was trained with local context of size $512$ and different $d$. FoT $2$->$64$ started with $d=2$ and then switched to $d=64$. Single-doc MT was trained with a memory size of $16K$. As we increase the memory size, the number of distractions (irrelevant keys) increases, making the task harder. Single-doc MT evaluated in the single-doc setting is a soft lower bound since it lacks distractions.
Appendix E Hyperparameters
Table 9 shows hyperparameters used in our experiments. We used context length $512$ unless stated otherwise. In Appendix F, Section 5.3, Section 5.5, we use the total batch size of $32K$ tokens. In Section 5.2 and Section 5.4, the total batch size is $128K$ tokens.
@@ Appendix E Hyperparameters
#Params $37M$ $184M$
Max learning rate $0.02$ $0.01$
Min learning rate $0.01$ $0.0005$
Embedding dim $512$ $1024$
Head dim $64$ $128$
#Heads $8$ $8$
FeedForward dim $2048$ $4096$
Local context length $256$/$512$ $512$
Batch size $32K$/$64K$ $32K$
#Number of training steps $5k$ $500k$
@@
For the dictionary lookup task (Appendix F) we checked the following hyperparameters:
@@
E.1 Schedule of $d$
In Sections 5.3, 5.4 and 5.5 for models with $d\in\{1,2,4,8\}$ we used constant schedule, and for models with $d=64$ we trained with $d=2$ for $450k$ steps and switched to $d=64$ for the final $50k$ steps. In Appendix F we trained with $d=1$ until the model reached $98\%$ accuracy and then we switched to $d=128$. For the $184M$ model in Section 5.2, we randomly sampled $d$ from $\{2,128\}$ in each training step.
Appendix F Dictionary lookup task
[…]
@@ Appendix F Dictionary lookup task
During training, we use documents comprising $512$ tokens. The first half of each document consists of definitions, whereas the second one consists of questions. For FoT, we use a local context of $256$, thus the model needs to use the memory attention layer to answer the questions correctly. We start with $d=1$ and increase to $d=128$ as soon as the model is able to reach $98\%$ training accuracy. During the inference, we use $k=32$ (the number of keys retrieved by kNN). As a baseline, we use a standard transformer model trained with the context length of $512$. In evaluation, we test different local context lengths, which quickly leads to very poor results.
In evaluation, we use longer documents but make only the last $256$ tokens correspond to questions. That is, as the context gets bigger (the token axis on Figure 10), the number of definitions increases, but the number of queries remains the same.
Appendix G FoT fine-tuning
For comparison in Table 3, our model is pre-trained for $100k$ steps with a total batch size of 128 ($128K$ tokens per step, with $1024$ local context). Then we fine-tune both FoT and baselines for additional $10k$ steps with the same batch size. When fine-tuning FoT, we randomly sample $d$ from $\{2,128\}$ in each training step to prevent the model from overfitting to a large additional context length during training.
Appendix H Datasets
GitHub
Isabelle
Appendix I Hardware and technical details
Appendix J Randomness
To evaluate the significance of our results, we conducted multiple runs for selected experiments in our study. In Figure 10, we calculate error bars, showing the minimum and maximum value over 10 runs of the same experiment. For the arXiv baseline experiment in Appendix K, we performed three runs with different random seeds and calculated their standard deviation, which is equal to $0.002$ perplexity. However, due to resource constraints, we were unable to conduct multiple runs for all experiments. Our preliminary findings indicate that the observed variance was minimal compared to the impact observed from other factors under investigation.
For the calculation of test perplexities, we used $1M$ tokens.
Appendix K Additional experimental results
Using the same setup, we study the performance of FoT while varying $d$ and $w$ configurations, similarly to Section 5.4, see Table 11. Parameter values $w=1$ and $w=2$ correspond to additional context lengths of $2K$ and $4K$, respectively. In an apples-to-apples comparison to MT with $2K$ additional context length, FoT outperforms the MT baseline, which shows the importance of trainable keys and values (see also Section C.1). Moreover, we confirm the findings from Section 5.4 that $d=2$ works significantly better than $d=1$ in all settings. Our best configuration achieves $2.148$ perplexity with $4K$ additional context during training, compared to $2.164$ of MT with $16K$ additional context.
Eval Context Training Memory Length
2k 8k 16k 32k
$4K$ 2.309 2.334 2.348 2.365
$8K$ 2.242 2.244 2.252 2.265
$16K$ 2.215 2.206 2.206 2.215
$32K$ 2.199 2.178 2.177 2.181
$64K$ 2.195 2.169 2.166 2.168
$128K$ 2.195 2.168 2.164 2.166
(1, 1), 2048 (1, 2), 2048 (2, 1), 4096 (2, 2), 4096
$4K$ 2.292 2.299 2.305 2.309
$8K$ 2.229 2.224 2.214 2.217
$16K$ 2.206 2.194 2.178 2.178
$32K$ 2.192 2.176 2.159 2.156
$64K$ 2.187 2.171 2.152 2.149
$128K$ 2.187 2.171 2.152 2.148
Appendix L Code
In Listing 2, we show the FoTs attention code (i.e., the code for the memory attention layers and crossbatch training), see Section 3, Appendix B.2, Appendix B.3. We note that the changes to the code are small; they are localized to the memory layer (the other layers follow the standard transformer protocol) and do not require any new trainable parameters.
Listing 2: Memory attention: Let $\ell$ be a memory attention layer. During the training, we make $\ell$ attend to the $(key,value)$ pairs from the local context, previous local context, and $d-1$ contexts coming from other documents. During the inference, queries from $\ell$ attend to the local context and $k$ nearest neighbors retrieved from memory. For simplicity, we provide the code for one head and assume that the crossbatch dimension $d$ is equal to the batch size.